Four mass–spring systems oscillate in simple harmonic motion. Rank the periods of oscillation for the mass–spring systems from largest to smallest.

m = 2 kg , k = 2 N/m

m = 2 kg , k = 4 N/m

m = 4 kg , k = 2 N/m

m = 1 kg , k = 4 N/m
Answer #1

\(T=2 \cdot \pi \sqrt{\frac{m}{k}}\)

\(1 \cdot \mathrm{m}=4 \mathrm{~kg}, \mathrm{k}=2 \mathrm{~N} / \mathrm{m}\)

\(2 m=2 \mathrm{~kg}, \mathrm{k}=2 \mathrm{~N} / \mathrm{m}\)

3. \(m=2 \mathrm{~kg}, \mathrm{k}=4 \mathrm{~N} / \mathrm{m}\)

4. \(m=1 \mathrm{~kg}, \mathrm{k}=4 \mathrm{~N} / \mathrm{m}\)